How do i find the area of a 45°-45°-90° triangle with a 4-cm hypotenuse
Answer by stanbon(75887)

How do i find the area of a 45°-45°-90° triangle with a 4-cm hypotenuse
-----
45-45l means the two legs are equal in length; let their length be "x":
Equation:
x^2 + x^2 = 4^2
2x^2 = 16
x^2 = 8
x = 2sqrt(2)
------------
Area = (1/2)(leg)(leg)
Area = (1/2)(8)
Area = 4 sqrt cm
======================
Cheers,
Stan H.